# Floating orbits and energy extraction from magnetized Kerr black holes

### v1

@@ IV Results
where $\dot{E}_{b}$ is the binding energy of the system. When the flux at the horizon is negligible (or positive), the balance above shows immediately that the particle has to inspiral: since we always have $\dot{E}^{\infty}>0$, to obey the balance equation the system needs to bind more strongly. However, when $\dot{E}^{\rm H}+\dot{E}^{\infty}<0$, as is the case here, the opposite happens: all the energy radiated at infinity is extracted from the BH rather than from binding energy. The particle outspirals rather than inspiralling. In fact, given other dissipation mechanisms, the particle may float.
Simply put, orbits inside $r_{0}\approx 9M$ shrink as a result of radiation reaction, because the particle is losing energy; conversely, orbits outside $r_{0}\approx 9M$ widen, since the particle is gaining energy. These later orbits are called floating orbits, and they do not exist in the absence of magnetic field [50] (except if there are massive fundamental fields present [51]). By continuity, there is a particular value of orbital radius $r_{0}\approx 9M$ for which the energy coming out of the horizon equals the energy going to infinity. For this value of $r_{0}$, we have a static orbit; the particle remains in the same orbital radius, neither losing nor gaining energy.
Finally, a few words on anti-resonances, i.e., the inverted peaks around $r_{0}\approx 10M$ and $r_{0}\approx 20M$ in Figs. 3 and 4, respectively. These are single orbits for which the flux at infinity and at the horizon are orders of magnitude lower than in very close orbits. This effect occurs for orbits with $\ell+m$ odd, at slightly different values of $r_{0}$ depending on $\ell,m$. Thus, this effect cannot be exclusively governed by the mode frequency $\omega=m\Omega_{0}$: otherwise, modes sharing the same azimuthal number would have anti-resonances at the same orbit. We have no convincing explanation for this effect, but it is worth noticing that the radiation in these modes is nevertheless suppressed as compared to the dominant the $\ell=m=1$ mode.
@@ V Conclusions
Magnetic fields are ubiquitous in nature. In this work, we studied radiation from charged particles around weakly magnetized BHs. We solved the relevant Teukolsy equation describing the EM radiation properties, both numerically and analytically (in the low-frequency regime). Both approaches agree very well in the overlapping regime of validity. Perhaps unsurprisingly, the dipolar mode dominates radiation emission. More interestingly, this radiation falls within the superradiant regime, that is, the energy flux at the horizon is negative. For $r_{0}\gtrapprox 9M$ the energy coming out of the horizon is larger than that lost at infinity: the particle extracts energy from the BH and will spiral outward, the first example of floating with standard model fundamental fields. Orbits satisfying this condition are called floating orbits and they do not occur the unmagnetized Kerr BH case.
We expect floating to occur also for gravitational and scalar radiation. This expectation is motivated by the fact that the “horizon dominance effect” is present in all three channels [49] and is the driver of the energy extraction mechanism. In astrophysical setups, charged particles will be part of an accretion disk or gas, and friction forces (i.e., nearest neighbhour interactions) are expected to dominate [23], leading to a net energy loss and thus destroying floating. Finally, the extension of our results to non-circular motion is both interesting and important, since it would test the robustness of the energy extraction mechanism. We leave this for future work.
@@ Acknowledgements.
We thank Rodrigo Vicente and Diogo Ribeiro for insightful discussions. This work makes use of the Black Hole Perturbation Toolkit. We acknowledge support by VILLUM Foundation (grant no. VIL37766) and the DNRF Chair program (grant no. DNRF162) by the Danish National Research Foundation. V.C. is a Villum Investigator and a DNRF Chair. J.S.S. and V.C. acknowledge financial support provided under the European Union’s H2020 ERC Advanced Grant “Black holes: gravitational engines of discovery” grant agreement no. Gravitas–101052587. Views and opinions expressed are however those of the author only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. J.N. was partially supported by FCT/Portugal through CAMGSD, IST-ID, projects UIDB/04459/2020 and UIDP/04459/2020. a the This project has received funding from the European Union’s Horizon 2020 research and innovation programme H2020-MSCA-2022-SE project EinsteinWaves, GA no. 101131233. and GRU, GA no. 101007855.
@@ Appendix A Circular orbit evolution
In this appendix we show that a charged particle, initially in a circular orbit in the equatorial plane, will radiate energy $E$ and angular momentum $L$ in precisely the correct ratio for it to remain in (quasi-)circular motion. This statement is proved assuming the adiabatic approximation is valid. As pointed out in Sec. III.3, this does not imply that the circular orbit is stable under radiation.
@@
where all the metric components are functions of $r$ and $\theta$ only. We also allow for an EM potential satisfying the same symmetry, $A_{\mu}=A_{\mu}(r,\theta)$. This symmetry is encoded in the existence of two Killing vector fields, $X^{\mu}$ and $Y^{\mu}$ (see Eq. (5)), that commute with the 4-potential. The particle’s energy $\mathcal{E}$ and angular momentum $\mathcal{L}$ per unit mass of the particle are still given by Eq. (9). We focus on the case where the motion is confined to the equatorial plane, which means that the values of $\mathcal{E}$ and $\mathcal{L}$ completely characterize the trajectory.
The Hamiltonian description of this system is identical to the one given in the main text, except that now we allow for $g_{\mu\nu}$ and $A_{\mu}$ to be generic stationary axisymmetric metrics and potentials. This means that Eq. (8) is not changed, but we rewrite $\mathcal{F}$ as
@@
$\bm{\Xi}^{T}$ $\cdot\mathcal{M}_{g}\cdot\bm{\Xi}=-1$ (40)
@@
where primes denote derivatives in $r$. Suppose the adiabatic approximation is valid, so that the particle is moving in a sequence of circular orbits described by $\mathcal{E}$ and $\mathcal{L}$, which are taken to be slowly varying functions of $r$. This is what we call quasicircular motion. Differentiating Eq. (40) gives

### v2

@@ IV Results
where $E_{b}$ is the binding energy of the system. When the flux at the horizon is negligible (or positive), the balance above shows immediately that the particle has to inspiral: since we always have $\dot{E}^{\infty}>0$, to obey the balance equation the system needs to bind more strongly. However, when $\dot{E}^{\rm H}+\dot{E}^{\infty}<0$, as is the case here, the opposite happens: all the energy radiated at infinity is extracted from the BH rather than from binding energy. The particle outspirals rather than inspiralling. In fact, given other dissipation mechanisms, the particle may float.
Simply put, orbits inside $r_{0}\approx 9M$ shrink as a result of radiation reaction, because the particle is losing energy; conversely, orbits outside $r_{0}\approx 9M$ widen, since the particle is gaining energy. These later orbits are called floating orbits, and they do not exist in the absence of magnetic field [51] (except if there are massive fundamental fields present [52]). By continuity, there is a particular value of orbital radius $r_{0}\approx 9M$ for which the energy coming out of the horizon equals the energy going to infinity. For this value of $r_{0}$, we have a stationary orbit; the particle remains in the same orbital radius, neither losing nor gaining energy.
To better visualize the floating and outspiraling effects mentioned above, we represent in Fig. 6 some orbital evolutions, using plots of the orbital radius $r_{0}$ as a function of coordinate time $t$. Concretely, we consider a charged particle with $q=1\,C$ and $m\approx 3.33\times 10^{-4}\,kg$, moving in the vicinity of a black hole with $M=10M_{\odot}$, for various cyclotron frequencies $\omega_{c}$. The value of $m$ is chosen so that $M\omega_{c}=0.1\iff B_{0}=10^{8}$ Gauss. The particle starts at $r_{0}=9M$, which is barely outside (inside) the floating orbit for $M\omega_{c}=0.1$, corresponding to the blue curve ($M\omega_{c}=0.09$, corresponding to the red curve). Thus, for $M\omega_{c}\leq 0.09$ we see an inspiral, while the for $M\omega_{c}\geq 0.1$ we see an outspiral. Notice how for the curves where $r_{0}=9M$ is close to being a floating orbit (blue and red) the particle lingers around this value of $r_{0}$ before spiraling in or out. Finally, the larger $\omega_{c}$ the smaller the orbital velocity (see Fig. 2), so the evolution happens on larger timescales. This is why the blue and green curves intersect each other.
Finally, a few words on anti-resonances, i.e., the inverted peaks around $r_{0}\approx 10M$ and $r_{0}\approx 20M$ in Figs. 3 and 4, respectively. These are single orbits for which the flux at infinity and at the horizon are orders of magnitude lower than for nearby orbits. This effect occurs for orbits with $\ell+m$ odd, at slightly different values of $r_{0}$ depending on $\ell,m$. Thus, this effect cannot be exclusively governed by the mode frequency $\omega=m\Omega_{0}$: otherwise, modes sharing the same azimuthal number would have anti-resonances at the same orbit. We have no convincing explanation for this effect, but it is worth noticing that the radiation in these modes is nevertheless suppressed when compared with the dominant the $\ell=m=1$ mode.
@@ V Conclusions
Magnetic fields are ubiquitous in nature. In this work, we studied radiation from charged particles around weakly magnetized BHs. We solved the relevant Teukolsy equation describing the EM radiation properties, both numerically and analytically (in the low-frequency regime). Both approaches agree very well in the overlapping regime of validity. Perhaps unsurprisingly, the dipolar mode dominates radiation emission. More interestingly, this radiation falls within the superradiant regime, that is, the energy flux at the horizon is negative. For $r_{0}\gtrapprox 9M$, the energy coming out of the horizon is larger than that lost at infinity: the particle extracts energy from the BH and will spiral outwards, the first example of floating with standard model fundamental fields. Orbits satisfying this condition are called floating orbits, and they do not occur in the unmagnetized Kerr BH case.
We expect floating to also occur for gravitational and scalar radiation. This expectation is motivated by the fact that the “horizon dominance effect” is present in all three channels [49], and is the driver of the energy extraction mechanism. In astrophysical setups, charged particles will be part of an accretion disk or ionized gas, and friction forces (i.e., nearest neighbhour interactions) are expected to dominate [23], leading to a net energy loss (thus destroying floating). Finally, the extension of our results to non-circular motion is both interesting and important, since it would test the robustness of the energy extraction mechanism. We leave this for future work.
@@ Acknowledgements.
We thank Rodrigo Vicente, Diogo Ribeiro and Martin Kološ for insightful discussions. This work makes use of the Black Hole Perturbation Toolkit. We acknowledge support by the VILLUM Foundation (grant no. VIL37766) and the DNRF Chair program (grant no. DNRF162) by the Danish National Research Foundation. V.C. is a Villum Investigator and a DNRF Chair. J.S.S. and V.C. acknowledge financial support provided under the European Union’s H2020 ERC Advanced Grant “Black holes: gravitational engines of discovery” grant agreement no. Gravitas–101052587. Views and opinions expressed are however those of the author only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. J.N. was partially supported by FCT/Portugal through CAMGSD, IST-ID, projects UIDB/04459/2020 and UIDP/04459/2020. This work has received funding from the European Union’s Horizon 2020 research and innovation programme H2020-MSCA-2022-SE through projects EinsteinWaves, GA no. 101131233, and GRU, GA no. 101007855.
@@ Appendix A Circular orbit evolution
In this appendix we show that a charged particle, initially in a circular orbit in the equatorial plane, will radiate energy $E$ and angular momentum $L$ in precisely the correct ratio for it to remain in (quasi-)circular motion. This statement is proved assuming that the adiabatic approximation is valid. As pointed out in Sec. III.3, this does not imply that the circular orbit is stable under radiation.
@@
where all the metric components are functions of $r$ and $\theta$ only. We also allow for an EM potential satisfying the same symmetry, $A_{\mu}=A_{\mu}(r,\theta)$. This symmetry is encoded in the existence of two Killing vector fields, $X^{\mu}$ and $Y^{\mu}$ (see Eq. (5)), that commute with the 4-potential. The particle’s energy per unit mass $\mathcal{E}$ and angular momentum per unit mass $\mathcal{L}$ are still given by Eq. (9). We focus on the case where the motion is confined to the equatorial plane, which means that the values of $\mathcal{E}$ and $\mathcal{L}$ completely characterize the trajectory.
The Hamiltonian description of this system is identical to that given in the main text, except that now we allow for $g_{\mu\nu}$ and $A_{\mu}$ to be generic stationary axisymmetric metrics and potentials. This means that Eq. (8) is not changed, but we rewrite $\mathcal{F}$ as
@@
$\bm{\Xi}^{T}$ $\cdot\mathcal{M}_{g}\cdot\bm{\Xi}=-1\,,$ (40)
@@
where primes denote derivatives in $r$. Suppose that the adiabatic approximation is valid, so that the particle is moving in a sequence of circular orbits described by $\mathcal{E}$ and $\mathcal{L}$, which are taken to be slowly varying functions of $r$. This is what we call quasicircular motion. Differentiating Eq. (40) gives